### v1

@@ 5 Proof for $1-\delta$ Quantile
In this section, we provide the calculation of the probability distribution of random variable $W:=\sum_{j=1}^{m}\mathbbm{1}\{g[j]\neq g^{\prime}[j]\}$, which plays an important part in the proof of the privacy guarantee for our algorithm (see Section 6). In Section 5.1, we present the definition of random variables $W,Y,Z$ used in this section. In Section 5.2, we calculate the probability distribution of $Y$. In Section 5.3, we calculate the probability distribution of $Z$ conditioned on $Y$. In Section 5.4, we calculate the probability distribution of $W$.
5.1 Definition
@@
Follow the Definition 5.2, let $Y_{x}$ denotes the set of distinct values in $y_{1},y_{2},\cdots,y_{k}$, and $Y_{x^{\prime}}$ denotes the set of distinct values in $y_{1}^{\prime},y_{2}^{\prime},\cdots,y_{k}^{\prime}$.
@@ 5.2 Distribution of $Y$
  • Let $y_{1},y_{2},\cdots,y_{k}$ be defined in Definition 5.2.
  • Let $Y$ be defined as Definition 5.2.
@@
For the failure case, we have $\sum_{i=1}^{k-1}\Pr[Y=i]\cdot\binom{m-i}{y-i}$. The $\binom{m-i}{y-i}$ term is due to repeated counting for each $i\in[k-1]$, where we first fix $i$ positions and then randomly pick the other $y-i$ different positions in the total $m-i$ positions.
@@ 5.3 Distribution of $Z$ conditioned on $Y$
Lemma 5.5 (Probability of $Z$ conditioned on $Y_{x}$ and $Y_{x^{\prime}}$).
@@
  • Let $Y_{x},Y_{x^{\prime}},Z$ be defined as Definition 5.3.
  • Let $A_{n}^{m}$ denotes $n!/(n-m)!$.
@@
$\Pr[|Z|=z||Y_{x}|=a,|Y_{x^{\prime}}|=b]=\frac{A_{m}^{a}\cdot\binom{b}{t}\cdot A_{m-a}^{t}\cdot A_{a}^{b-t}}{A_{m}^{a}\cdot A_{m}^{b}}.$
@@
Since $Z$ is defined as Definition 5.3, we can have the following
@@
We also need to choose $t$ different positions in the rest $m-a$ positions for $|Y_{x^{\prime}}|-|Y_{x}\cap Y_{x^{\prime}}|$ part. Hence, we have the $A_{m-a}^{t}$ term in the numerator.
Lastly, let’s consider the $b-t$ part. For this part, we need to choose $b-t$ different positions from $a$ positions. Therefore, we have the $A_{a}^{b-t}$ term in the numerator.
@@
$\Pr[|Z|=z||Y_{x}|=a,|Y_{x^{\prime}}|=b]=\frac{A_{m}^{a}\cdot\binom{b}{t}\cdot A_{m-a}^{t}\cdot A_{a}^{b-t}}{A_{m}^{a}\cdot A_{m}^{b}}.$
5.4 Distribution of $W$
Figure 1: Let $W:=|S|$ denote the number of bits in the Bloom filter changed by substituting an element in the inserted set $A$ (Definition 3.2). We achieve $\epsilon_{0}$-DP for each single bit and $(\epsilon,\delta)$-DP for the entire Bloom filter via the random response (Definition 3.4), where $\epsilon_{0}=\epsilon/N$. The $N$ is $1-\delta$ quantile of the random variable $W$. We visualize the distribution of the random variable $W$ (see Lemma 5.6) under the setting described in the experiments section (Section 9). Namely, we have the bit array length in the Bloom filter $m=2^{19}$, the number of elements inserted into the Bloom filter $|A|=10^{5}$, and the number of hash functions $k=3$. It can be inferred from this visualization that the values of random variable $W$ have good concentration properties, mostly concentrated around its mean.
@@ 5.4 Distribution of $W$
  • Let $Y_{x},Y_{x^{\prime}},Z$ be defined as Definition 5.3.
  • Let $W$ be defined as Definition 5.1.
@@
  • Let $p_{0}:=(1-\frac{1}{m})^{(|A|-1)k}$ denotes the proportion of bits which are still 00 in the bit-array.
@@
We can get the probability of $\Pr[|Y_{x}|=a]$ and $\Pr[|Y_{x^{\prime}}|=b$ from Lemma 5.4.
We can get the probability of $\Pr[|Z|=z||Y_{x}|=a,|Y_{x^{\prime}}|=b]$ from Lemma 5.5.
Now, let’s consider the $\Pr[W=w||Z|=z,|Y_{x}|=a,|Y_{x^{\prime}}|=b]$ term.
Note that only elements in the exclusive-or set may contribute to the final $W$. Therefore, we have $w\leq n_{2}$. Namely, when $n_{2}<w$, we have $\Pr[W=w||Z|=z,|Y_{x}|=a,|Y_{x^{\prime}}|=b]=0$.
Now, let’s calculate $\Pr[W=w||Z|=z,|Y_{x}|=a,|Y_{x^{\prime}}|=b]$ under $n_{2}\geq w$ condition.
@@
  • We define $Z_{zero}$ as the set of positions of bits which are still 00 after the insertion of $A_{fix}$.
@@
6 Privacy guarantees for one coordinate
@@
In Section 6.1, we demonstrate the privacy guarantees for single bit of array in Bloom filter.
Then in Section 6.2, we provide the proof of privacy guarantees for our entire algorithm.
6.1 Single bit is private
@@
Lemma 6.1 (Single bit is private ).
@@
6.2 Privacy guarantees for DPBloomfilter
@@
Theorem 6.2 (Privacy for Query, formal version of Lemma 4.1).
@@
  • Let $N=F_{W}^{-1}(1-\delta)$ denote the $1-\delta$ quantile of the random variable $W$ (see Definition 5.1).
@@
By the definition of $N$, we know that with probability at least $1-\delta$, $|S|\leq F^{-1}(1-\delta)=N$. Hence, Eq. (6.2) is upper bounded by $\epsilon$ with probability $1-\delta$.
@@
7 Utility analysis
In this section, we establish the utility guarantees for our algorithm. Initially, we calculate the accuracy for the query of the standard Bloom filter in Section 7.1. We then assess the utility loss caused by introducing the random response technique by comparing the output of the DPBloomfilter with the output of the standard Bloom filter in Section 7.2. Ultimately, we present the assessment of our algorithm’s utility in Section 7.3.
@@ 7 Utility analysis
Definition 7.1.
Let $z\in\{0,1\}$ denote the true answer for whether $x\in A$. Let $\widehat{z}\in\{0,1\}$ denote the answer outputs by Bloom filter. Let $\widetilde{z}\in\{0,1\}$ denote the answer output by DPBloomFilter (Algorithm 1).
7.1 Accuracy for query of Standard Bloom Filter
@@
Lemma 7.2 (Accuracy for query of Standard Bloom Filter ).
@@
  • Let $\widehat{z}$ be defined as Definition 7.1.
  • Let $z$ be defined as Definition 7.1.
@@
If we have inserted $|A|$ elements, the probability that a certain bit is still 00 is
@@
where the first step follows from Bloom filter only has false positive error, the second step follows from the definition of event $E_{1}$ and the definition of $\alpha$, the third step follows from Eq. (7.1).
7.2 Accuracy (compare DPBloomFilter with Standard BloomFilter) for Query
@@
Lemma 7.3 (Accuracy (compare DPBloomFilter with Standard BloomFilter) for Query ).
@@
  • Let $\widehat{z}$ be defined as Definition 7.1.
  • Let $\widetilde{z}$ be defined as Definition 7.1.
@@
  • Let $\delta_{\mathrm{err}}$ be defined as in Lemma 7.2.
@@
$[\widetilde{z}=1|\widehat{z}=0]$ happens only if the following conditions hold:
@@
    All elements in $Z$ flip from 00 to $1$.
@@
$=$ $~{}\Pr[\widetilde{z}=0|\widehat{z}=0]\cdot\widehat{\alpha}+\Pr[\widetilde{z}=1|\widehat{z}=1]\cdot(1-\widehat{\alpha})$
@@
where the first step follows from the definition of $\widehat{\alpha}$, the second step follows from $\widehat{\alpha}=\alpha(1-\delta)$, the third step follows from Eq. (6) Eq. (7.2), the fourth step follows from basic algebra rules, the fifth step follows from $(1-\alpha+\alpha\cdot\delta_{\mathrm{err}})\geq 0$.
@@
7.3 Accuracy (compare DPBloomfilter with true-answer) for Query
@@
Theorem 7.4 (Accuracy (compare DPBloomfilter with true-answer) for Query, formal version of Lemma 4.2).
@@
  • Let $\widehat{z}$ be defined as Definition 7.1.
  • Let $z$ be defined as Definition 7.1.
@@
  • Let $\delta_{\mathrm{err}}$ be defined as in Lemma 7.2.
@@
where the first step from basic probability rules, the secod step follows from Equation 3, Equation 7.2 and definition of $\alpha$ and $t$, the third step follows from basic algebra, the fourth step follows from the fact that $t,\alpha\in[0,1]$.
@@
Lemma 7.5 (Accuracy for Query under Random Guess).
@@
  • Let $\widehat{z}$ be defined as Definition 7.1.
  • $\epsilon_{0}=0$. Namely, each bit in the bit-array of the DP Bloom has $\frac{1}{2}$ probability to be set to 00, and $\frac{1}{2}$ probability to be set to $1$.
@@
8 Running Time
In this section, we provide the proof of running time for Algorithm 1. The running time for our algorithm consists of two parts: time for initialization in Section 8.1 and time for query in Section 8.2.
8.1 Running time for initialization
@@
Lemma 8.1 (Running time for initialization).
@@
8.2 Running time for query
@@
Lemma 8.2 (Running time for query).
@@
By combing the result of Lemma 8.1 and Lemma 8.2, we can obtain the running of our entire algorithm is $O(|A|\cdot k\cdot\mathcal{T}_{h}+m)$.
9 Experiments
Figure 2: Three kinds of error rates with different bit-array lengths $m$. We fix the number of inserted elements $|A|=10^{5}$, the number of hash functions $k=3$, and $\delta=0.01$ in $(\epsilon,\delta)$-DP. In the figure, $\log$ denotes $\log_{2}$. Left: Total error denotes the case when we randomly choose queries from the universe $[n]$; Middle: False negative denotes the case when we randomly choose queries from the set $S$, which represents the set of elements inserted into the DP Bloom filter; Right: False positive denotes the case when we randomly choose queries from the set $\overline{S}=[n]\backslash S$. As $m$ increases, the total error rate and false positive error rate decrease accordingly, while false negative error rate remains constant. As $\epsilon$ approaches 00, the DP Bloom filter gets closer to random guessing. In this case, the false positive error rate converges to $\frac{1}{2^{k}}$, and the false negative error rate converges to $1-\frac{1}{2^{k}}$. This is consistent with our result in Lemma 7.5 Our DPBloomFilter achieves practical utility when $\epsilon$ is small(e.g. $\epsilon<10$).
Figure 3: Three kinds of error rates with different numbers of inserted elements $|A|$. We fix the length of bit-array $m=2^{19}$, the number of hash functions $k=3$, and $\delta=0.01$ in $(\epsilon,\delta)$-DP. As $|A|$ increases, the Total Error Rate and false positive error rate increase accordingly, while the false negative error rate remains constant.
Figure 4: Three kinds of error rates with different numbers of hash function $k$. We fix the length of bit-array $m=2^{19}$, the number of inserted elements $|A|=10^{5}$, and $\delta=0.01$ in $(\epsilon,\delta)$-DP. As $k$ increases, the Total Error Rate and false positive error rate decrease accordingly, while the false negative error rate increases accordingly.
In this section, we introduce the simulation experiments conducted on the DPBloomfilter. In Section 9.1, we introduce the basic setup of our experiments and restate basic definitions of three kinds of error. In Section 9.2, we discuss the results of our experiments, which align with our theoretical analysis.
9.1 Experiments Setup and Basic Notations
Recall that we have the following notations. Let $m$ denote the length of the bit array in the DPBloomfilter. Let $|A|$ denote the number of elements inserted into the DPBloomfilter. Let $k$ denote the number of hash functions used in the DPBloomfilter. Let $\epsilon,\delta$ denote the differential privacy parameters of the DPBloomfilter. Let $N$ denotes the $1-\delta$ quantile of $W$ (see Definition 5.1), and the close-form of the distribution of $W$ is shown in Lemma 5.6. Let $\epsilon_{0}=\epsilon/N$. By Theorem 4.1, we choose $\epsilon_{0}$ in this way can guarantee to $(\epsilon,\delta)$-DP in the whole algorithm. Unless specified, we adopt $m=2^{19},|A|=10^{5},k=8,n=2^{63}\approx 10^{19}$ in the following experiments. We choose this $n$ because this $n$ is the biggest integer that can be represented on our server.
Recall that $[n]$ denotes the universe. Let $S$ denote the elements inserted into the DPBloomfilter. Let $\overline{S}=[n]\backslash S$ denote the elements not inserted into the DPBloomfilter. Let $\widetilde{z}\in\{0,1\}$ denote the answer output by DPBloomfilter.
We report three kinds of error rates in our experiments. They are the following: (1) total error, where we randomly choose queries from the universe $[n]$ and report the error rate of our DPBloomfilter; (2) false positive error, where we random choose queries from $\overline{S}$. When the DPBloomfilter outputs $\widetilde{z}=1$, this will cause a false positive error; (3) false negative error, where we random choose queries from $S$. When the DPBloomfilter outputs $\widetilde{z}=0$, this will cause a false negative error.
9.2 Experiment Results
In this section, we conduct experiments based on the setting mentioned in the previous section. Specifically, we run simulation experiments on different $m$, $|A|$, and $k$ to demonstrate the utility of our algorithm under differential privacy guarantees.
In Figure 2, we conduct experiments on different $m$, whereas $m$ increases, the total error rate and false positive error rate decrease accordingly, while the false negative error rate remains constant.
In Figure 3, we also conduct experiments on different $|A|$, whereas $|A|$ increases, the total error rate and false positive error rate increase accordingly. At the same time, the false negative error rate remains constant. This phenomenon is consistent with our theoretical analysis of the utility of DPBloomfilter (Theorem 4.2). Recall that we have $\alpha=\Pr[z=0]$, denoting the probability of an arbitrary query $q\notin A$. Since $|A|$ increases, $\alpha$ decreases, the utility guarantee in Theorem 4.2, which is consistent with higher error rate in our experiment results.
In Figure 4, we conduct experiments on different $k$ as well, whereas $k$ increases, the total error rate, and false positive error rate decrease, while the false negative error rate increases accordingly.
Note that in Figure 2, Figure 3, and Figure 4, as $\epsilon$ approaches 00, the DPBloomfilter gets closer to random guessing. In this case, the false positive error rate converges to $\frac{1}{2^{k}}$, and the false negative error rate converges to $1-\frac{1}{2^{k}}$. This is consistent with our result in Lemma 7.5. Also, as $\epsilon$ increases, the three types of error rates in the Bloom filter with differential privacy (DP) approach the error rates observed when DP is not applied. This is consistent with the intuition that when $\epsilon$ increases, there is less privacy. Therefore, the performance approaches the performance of a Bloom filter without any privacy guarantees.
10 Discussion
Section 10.1 discusses why the random response mechanism is preferred over Gaussian and Laplace mechanisms for achieving differential privacy. In Section 10.2, we consider the underlying reasons for applying the random response mechanism to both $1$ and 00.
10.1 Why Random Response but not Gaussian or Laplace Noise?
As mentioned in Section 2, Gaussian and Laplace noise are two classical mechanisms to achieve differential privacy.
The advantage of the Laplace mechanism is that its distribution is concentrated on its mean. Under the same privacy budget, it will not introduce too much noise like the Gaussian mechanism due to the long-tail nature of its distribution. The advantage of the Gaussian mechanism is that it has good mathematical properties and makes it easy to analyze the utility of private data structures.
However, the above two mechanisms are not as effective as the random response (flip coin) mechanism when dealing with discrete values. Here, we consider the case where the discrete values are integers. Under certain privacy budgets, the noise added by Gaussian and Laplace mechanisms does not reach the threshold of $0.5$, resulting in attackers being able to remove the noise through rounding operations easily, and the privacy of the data structure no longer exists.
In our case, each bit of the Bloom filter can only be $1$ or 00, which is consistent with the above situation. Hence, our work only considers the random response mechanism instead of classical Gaussian and Laplace mechanisms.
10.2 Why Flip Both 00 and $1$?
In our work, we apply random response mechanism to each bit in the Bloom filter, either it is 00 or $1$. Although this will lead to a certain probability of false negatives in the Bloom filter, we argue that it is necessary to make the Bloom filter differentially private.
Let’s consider what will happen if we don’t apply random response mechanism like this. Suppose we only apply random responses to bits that are $1$ in the Bloom filter and leave the bits with 0 untouched. Following the notations used in Lemma A, we use $g\in\{0,1\}^{m}$ to represent the bit array generated by inserting the original dataset into the Bloom filter and $g^{\prime}\in\{0,1\}^{m}$ to represent the bit array generated by inserting the neighboring dataset into the Bloom filter. We use $\widetilde{g}$ and $\widetilde{g}^{\prime}$ to denote their private version, respectively. Without loss of generality, for some $j\in[m]$, we assume $g[j]=1$ and $g^{\prime}[j]=0$. Since we only apply random response mechanism on bits with value $1$, then $\Pr[\widetilde{g}^{\prime}[j]=1]=0$. Therefore, we cannot calculate $\Pr[\widetilde{g}[j]=1]/\Pr[\widetilde{g}^{\prime}[j]=1]$, since the denominator is 00. Hence, we cannot have any privacy guarantees under this setting. Similar situations occur when we apply a random response mechanism on bits with value 00. We also cannot prove the differential privacy property of the Bloom filter. Therefore, we have to apply the random response mechanism on bits either with value 00 or $1$.
11 Conclusion
In this work, we propose DPBloomfilter, a novel method ensuring the privacy of the Bloom filter via random response. To the best of our knowledge, this is the first work applying the random response mechanism to achieve DP on the membership query task of the Bloom filter. From the privacy side, we have proved that our method achieves $(\epsilon,\delta)$-DP with the same running complexity as the standard Bloom filter. From the utility side, we demonstrate from both theoretical and experimental perspectives that the Bloom filter can ensure high utility while ensuring differential privacy.
Acknowledgement
Research is partially supported by the National Science Foundation (NSF) Grants 2023239-DMS, CCF-2046710, and Air Force Grant FA9550-18-1-0166.

### v2

@@ Appendix B Proof for $1-\delta$ Quantile
In this section, we provide the calculation of the probability distribution of random variable $W:=\sum_{j=1}^{m}\mathbbm{1}\{g[j]\neq g^{\prime}[j]\}$, which plays an important part in the proof of the privacy guarantee for our algorithm (see Section C). In Section B.1, we present the definition of random variables $W,Y,Z$ used in this section. In Section B.2, we calculate the probability distribution of $Y$. In Section B.3, we calculate the probability distribution of $Z$ conditioned on $Y$. In Section B.4, we calculate the probability distribution of $W$.
B.1 Definition
@@
Follow the Definition B.2, let $Y_{x}$ denotes the set of distinct values in $y_{1},y_{2},\cdots,y_{k}$, and $Y_{x^{\prime}}$ denotes the set of distinct values in $y_{1}^{\prime},y_{2}^{\prime},\cdots,y_{k}^{\prime}$.
@@ B.2 Distribution of $Y$
  • Let $y_{1},y_{2},\cdots,y_{k}$ be defined in Definition B.2.
  • Let $Y$ be defined as Definition B.2.
@@
For the failure case, we have $\sum_{i=1}^{k-1}\Pr[Y=i]\cdot\binom{m-i}{y-i}$. The $\binom{m-i}{y-i}$ term is due to repeated counting for each $i\in[k-1]$, where we first fix $i$ positions and then randomly pick the other yiy-i different positions in the total mim-i positions.
@@ B.3 Distribution of $Z$ conditioned on $Y$
Lemma B.5 (Probability of $Z$ conditioned on $Y_{x}$ and $Y_{x^{\prime}}$).
@@
  • Let $Y_{x},Y_{x^{\prime}},Z$ be defined as Definition B.3.
  • Let $A_{n}^{m}$ denotes n!/(nm)!n!/(n-m)!.
@@
[…]
@@
Since $Z$ is defined as Definition B.3, we can have the following
@@
We also need to choose $t$ different positions in the rest mam-a positions for $|Y_{x^{\prime}}|-|Y_{x}\cap Y_{x^{\prime}}|$ part. Hence, we have the $A_{m-a}^{t}$ term in the numerator.
Lastly, let’s consider the btb-t part. For this part, we need to choose btb-t different positions from $a$ positions. Therefore, we have the $A_{a}^{b-t}$ term in the numerator.
@@
[…]
B.4 Distribution of $W$
Figure 1: Let $W:=|S|$ denote the number of bits in the Bloom filter changed by substituting an element in the inserted set AA (Definition 3.2). We achieve $\epsilon_{0}$-DP for each single bit and $(\epsilon,\delta)$-DP for the entire Bloom filter via the random response (Definition 3.4), where $\epsilon_{0}=\epsilon/N$. The NN is $1-\delta$ quantile of the random variable $W$. It can be inferred from this visualization that the values of random variable $W$ have good concentration properties, mostly concentrated around its mean.
@@ B.4 Distribution of $W$
  • Let $Y_{x},Y_{x^{\prime}},Z$ be defined as Definition B.3.
  • Let $W$ be defined as Definition B.1.
@@
  • Let $p_{0}:=(1-\frac{1}{m})^{(|A|-1)k}$ denotes the proportion of bits which are still 0 in the bit-array.
@@
We can get the probability of $\Pr[|Y_{x}|=a]$ and $\Pr[|Y_{x^{\prime}}|=b$ from Lemma B.4.
[…]
@@
  • We define $Z_{zero}$ as the set of positions of bits which are still 0 after the insertion of $A_{fix}$.
@@
Appendix C Privacy guarantees for one coordinate
@@
In Section C.1, we demonstrate the privacy guarantees for single bit of array in Bloom filter.
Then in Section C.2, we provide the proof of privacy guarantees for our entire algorithm.
C.1 Single bit is private
@@
Lemma C.1 (Single bit is private ).
@@
C.2 Privacy guarantees for DPBloomfilter
@@
Theorem C.2 (Privacy for Query, formal version of Lemma 4.1).
@@
  • Let $N=F_{W}^{-1}(1-\delta)$ denote the $1-\delta$ quantile of the random variable $W$ (see Definition B.1).
@@
By the definition of $N$, we know that with probability at least $1-\delta$, $|S|\leq F^{-1}(1-\delta)=N$. Hence, Eq. (C.2) is upper bounded by $\epsilon$ with probability $1-\delta$.
@@
Appendix D Utility analysis
In this section, we establish the utility guarantees for our algorithm. Initially, we calculate the accuracy for the query of the standard Bloom filter in Section D.1. We then assess the utility loss caused by introducing the random response technique by comparing the output of the DPBloomfilter with the output of the standard Bloom filter in Section D.2. Ultimately, we present the assessment of our algorithm’s utility in Section D.3.
@@ Appendix D Utility analysis
Definition D.1.
[…]
D.1 Accuracy for query of Standard Bloom Filter
@@
Lemma D.2 (Accuracy for query of Standard Bloom Filter ).
@@
  • Let $\widehat{z}$ be defined as Definition D.1.
  • Let $z$ be defined as Definition D.1.
@@
If we have inserted $|A|$ elements, the probability that a certain bit is still 0 is
@@
where the first step follows from Bloom filter only has false positive error, the second step follows from the definition of event $E_{1}$ and the definition of $\alpha$, the third step follows from Eq. (D.1).
D.2 Accuracy (compare DPBloomFilter with Standard BloomFilter) for Query
@@
Lemma D.3 (Accuracy (compare DPBloomFilter with Standard BloomFilter) for Query ).
@@
  • Let $\widehat{z}$ be defined as Definition D.1.
  • Let $\widetilde{z}$ be defined as Definition D.1.
@@
  • Let $\delta_{\mathrm{err}}$ be defined as in Lemma D.2.
@@
$\widetilde{z}=1|\widehat{z}=0]$ happens only if the following conditions hold:
@@
    All elements in $Z$ flip from 0 to $1$.
@@
[…]
@@
where the first step follows from the definition of $\widehat{\alpha}$, the second step follows from $\widehat{\alpha}=\alpha(1-\delta)$, the third step follows from Eq. (7) Eq. (D.2), the fourth step follows from basic algebra rules, the fifth step follows from $(1-\alpha+\alpha\cdot\delta_{\mathrm{err}})\geq 0$.
@@
D.3 Accuracy (compare DPBloomfilter with true-answer) for Query
@@
Theorem D.4 (Accuracy (compare DPBloomfilter with true-answer) for Query, formal version of Lemma 4.2).
@@
  • Let $\widehat{z}$ be defined as Definition D.1.
  • Let $z$ be defined as Definition D.1.
@@
  • Let $\delta_{\mathrm{err}}$ be defined as in Lemma D.2.
@@
where the first step from basic probability rules, the secod step follows from Equation 4, Equation D.2 and definition of $\alpha$ and $t$, the third step follows from basic algebra, the fourth step follows from the fact that $t,\alpha\in[0,1]$.
@@
Lemma D.5 (Accuracy for Query under Random Guess).
@@
  • Let $\widehat{z}$ be defined as Definition D.1.
  • $\epsilon_{0}=0$. Namely, each bit in the bit-array of the DP Bloom has $\frac{1}{2}$ probability to be set to 0, and $\frac{1}{2}$ probability to be set to $1$.
@@
Appendix E Running Time
In this section, we provide the proof of running time for Algorithm 1. The running time for our algorithm consists of two parts: time for initialization in Section E.1 and time for query in Section E.2.
E.1 Running time for initialization
@@
Lemma E.1 (Running time for initialization).
@@
E.2 Running time for query
@@
Lemma E.2 (Running time for query).
@@
By combing the result of Lemma E.1 and Lemma E.2, we can obtain the running of our entire algorithm is $O(|A|\cdot k\cdot\mathcal{T}_{h}+m)$.